# Information Propagation in Predator-Prey Dynamics of Turbulent Plasma

### v1

@@ Introduction.
Figure 2: (a) Schematic of information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$. (b) $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. The dashed line denotes the resonance condition $2\epsilon\beta-(1-\epsilon)^{2}=0$. The stochastic amplification occurs in the parameter region above this line.
@@
The information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$ quantifies the rate at which turbulence gains information about zonal flow. In particular, its sign indicates the direction in which the information is transmitted: if $\dot{I}^{Z\rightarrow W}_{\lambda}>0$ for all $\lambda$, the information of zonal flow is transferred to turbulence, and vice versa. Here, we note that the non-bipartite structure allows the case where the sign of $\dot{I}^{Z\rightarrow W}_{\lambda}$ varies depending on $\lambda$, which makes it difficult to determine the direction of information transfer [36]. In that case, $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ plays a special role because it is odd under time reversal and vanishes in the equilibrium state. Fortunately, as will be shown below, this problem is irrelevant to our model.
@@
where $p_{\tau}(z)$ and $p_{\tau}(w)$ are marginal distributions for zonal flow and turbulence, respectively. By noting that the steady-state distribution $p_{\mathrm{ss}}({\bm{r}})$ for the linear Langevin equations (5) and (6) is Gaussian with covariance matrix $\Sigma:=\langle({\bm{r}}-\langle{\bm{r}}\rangle)({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\rangle$, the information flow in the steady state can be calculated as
@@
where $\Sigma^{ZW}=\langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle$ and $\Sigma^{WW}=\langle(w-\langle w\rangle)^{2}\rangle$. See Ref. [47] for the detailed derivation of Eqs. (9) and (10). The inequality (10) states that the information about fluctuations of zonal flow is propagated to turbulence. In other words, the turbulence is “learning” about the fluctuating behavior of the zonal flow. This result is somewhat counterintuitive because the system exhibits persistent predator-prey cycles, where the fluctuations of zonal flow follow those of turbulence with a time lag. Figure 2(b) shows the $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. Interestingly, Fig. 2(b) suggests that the behavior of the information flow does not depend significantly on whether stochastic amplification occurs or not. In other words, even when the system does not exhibit the fluctuating predator-prey oscillation, there is a statistical causality from zonal flow to turbulence. This information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow. We finally remark that the inferred causality from the information flow is consistent with that from the cross-correlation function, although interpreting the latter is less straightforward [47].
Concluding remarks.
@@
—We have proposed a simple stochastic predator-prey model for plasma turbulence that exhibits quasi-cycles induced by stochastic amplification of intrinsic noise. For this model, we have quantified the causality between zonal flow and turbulence by showing that the information of zonal flow is propagated to turbulence.
@@
Our model illustrates the significance of considering the effects of intrinsic noise in understanding the predator-prey-like oscillations observed in plasma turbulence. Notably, our results suggest the possibility that the observed cyclic oscillations in a toroidal plasma are quasi-cycles rather than limit cycles. Using the methods proposed in Ref. [48], it would be possible to distinguish quasi-cycles and limit cycles both in experiments and simulations. Further studies are needed to verify whether both quasi-cycles and information transfer from zonal flow to turbulence can be observed in experiments and simulations of more realistic models, such as the Hasegawa–Wakatani equation [50].

### v2

@@ V Information propagation between zonal flow and turbulence
Figure 2: (a) Schematic of information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$. (b) $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ with $D^{W}=1/2$. The dashed line denotes the resonance condition $2\epsilon\beta-(1-\epsilon)^{2}=0$. The stochastic amplification occurs in the parameter region above this line.
@@
The information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$ quantifies the rate at which turbulence gains information about zonal flow. In particular, its sign indicates the direction in which the information is transmitted: If $\dot{I}^{Z\rightarrow W}_{\lambda}>0$ for all $\lambda$, the information of zonal flow is transferred to turbulence, and vice versa. Here, we note that the non-bipartite structure allows the case where the sign of $\dot{I}^{Z\rightarrow W}_{\lambda}$ varies depending on $\lambda$, which makes it difficult to determine the direction of information transfer [37]. In that case, $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ plays a special role because it is antisymmetric under time reversal and vanishes in the equilibrium state. Fortunately, as will be shown below, this problem is irrelevant to our model.
@@
where $p_{\tau}(z)$ and $p_{\tau}(w)$ are marginal distributions for zonal flow and turbulence, respectively. By noting that the steady-state distribution $p_{\mathrm{ss}}({\bm{r}})$ is Gaussian [Eq. (13)], the information flow in the steady state can be calculated as
@@
where $\Sigma^{ZW}=\langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle$ and $\Sigma^{WW}=\langle(w-\langle w\rangle)^{2}\rangle$. See Appendix B for the detailed derivation of Eqs. (20) and (21). The inequality (21) states that the information about fluctuations of zonal flow is propagated to turbulence. In other words, the turbulence is “learning” about the fluctuating behavior of the zonal flow. This result is somewhat counterintuitive because the system exhibits persistent predator-prey cycles, where the fluctuations of zonal flow follow those of turbulence with a time lag. Figure 2(b) shows the $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. Interestingly, Fig. 2(b) suggests that the behavior of the information flow does not depend significantly on whether stochastic amplification occurs or not. In other words, even when the system does not exhibit the fluctuating predator-prey oscillation, there is a predictive causality from zonal flow to turbulence. This information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow. We finally remark that the inferred causality from the information flow is consistent with that from the cross-correlation function, although interpreting the latter is less straightforward (see Appendix C).
VI Concluding remarks
@@
We have proposed a simple stochastic predator-prey model for plasma turbulence that exhibits quasi-cycles induced by stochastic amplification of intrinsic noise. For this model, we have quantified the predictive causality between zonal flow and turbulence by showing that the information of zonal flow is propagated to turbulence.
Several remarks on our results are in order. First, although we have considered the four simple “reactions” that incorporate a minimal self-regulation mechanism, it is straightforward to extend our approach to more complicated reactions, such as those involving time-dependent rate constants or additional reactant species (e.g., the mean shear flow). Even in such a case, both the system size expansion and the information-theoretic analysis are applicable.
@@
Second, our model illustrates the significance of considering the effects of intrinsic noise in understanding the predator-prey-like oscillations observed in plasma turbulence. Notably, our results suggest the possibility that the observed cyclic oscillations in a toroidal plasma are quasi-cycles rather than limit cycles. Using the methods proposed in Ref. [43], it would be possible to distinguish quasi-cycles and (stochastic) limit cycles both in experiments and simulations. Further studies are needed to verify whether both quasi-cycles and information transfer from zonal flow to turbulence can be observed in experiments and simulations of more realistic models, such as the Hasegawa–Wakatani equation [58].
Third, our results essentially hold true for general predator-prey dynamics, not limited to plasma turbulence. In particular, while the essence of our first main result—the emergence of quasi-cycles due to stochastic amplification in the predator-prey dynamics—has already been recognized in the field of ecology [22, 23], our second main result concerning information flow may offer new perspectives even within an ecological context.
Finally, while we have shown that information of zonal flow is propagated to turbulence, further research is needed to understand how this result relates to the regulation of turbulence through the control of zonal flow. This issue also seems connected to the general question of how predictive causality relates to dynamical causality, which can be quantified in frameworks such as convergent cross mapping [59, 60]. We hope our research will serve as a starting point for investigating these issues.